OpenStudy (anonymous):

what is the value of d when 11(d+1)=4(d+8)?????

OpenStudy (anonymous):

11d+11=4d+32 7d=21 d=3

OpenStudy (anonymous):

11(d+1)=4(d+8) distribute 11d+11=4d+32 subtract 4d from both sides, and subtract 11 from both sides 7d=21 divide both sides by 7 d = 3

OpenStudy (anonymous):

yeah its 3